# Forbidden Electron Transfer in the Adiabatic Limit of the Marcus-Inverted Region

### v2

@@ Abstract
Here it is shown that in the adiabatic limit of condensed-phase electron transfer, the onset of barrierless transition occurs at a lower driving force than predicted by the non-adiabatic Marcus formulation. Furthermore, in the adiabatic limit of the Marcus-inverted region, isoenergetic electron transfer is strictly forbidden in the absence of nuclear tunneling. This "forbidden" behavior arises from a topological change in the mapping between the adiabatic and diabatic electronic surfaces, emerging precisely at the onset of the Marcus-inverted region.
@@
but differences can be enormous when $V$ is not much smaller than $\lambda.$ For instance, for aqueous electroreduction of carbon dioxide on a gold electrode in the presence of potassium ions, the reorganization energy has been fitted experimentally by Zhang et al. as $\lambda_{\text{eff}}=0.75$ eV [4], while Qin et al. have measured $\lambda=6.3$ eV in ab initio simulations of the same system [3]. Using these values in Eq. (1) predicts $V=2.1$ eV, which has been corroborated by our own recent ab initio calculation of the coupling value [17]. These results may be critical to quantitatively understand the kinetics of strongly adiabatic electrochemical reactions.
@@
(iii) $\Delta G^{\circ}<-\lambda$: ET in this inverted region is forbidden and can only be achieved via nuclear tunneling, which is generally negligible compared to typical finite-temperature rates.
To justify these findings, we proceed with the analysis presented in Ref. [16], which represents the Marcus system in the basis of donor (D) and acceptor (A) diabatic states, parametrized by a solvent polarization coordinate $q$. The Hamiltonian is given by
@@
Figure 1: The two level Marcus system under study, where $q$ is the non-dimensionalized reaction coordinate and $E$ is the free energy. The system is shown in (a) the normal region (b) the transition from the normal region to the barrierless region, (c) the barrierless region, and (d) the forbidden region.
@@
$k_{\text{ET}}\approx\nu\exp\!\left\{-\frac{\alpha V^{2}}{\lambda}-\frac{(\lambda+\Delta G^{\circ})^{2}}{4\lambda k_{\mathrm{B}}T}\right\}.$ (10)
where $\nu$ is the classical attempt frequency. The adiabatic limit implies $\alpha V^{2}/\lambda\gg 1$, and so $k_{\text{ET}}\rightarrow 0$ regardless of the value of the Marcus activation factor. This result reveals a previously unrecognized property of the Marcus inverted region: in addition to the fact that increasing the thermodynamic bias suppresses the reaction rate, so too does increasing the coupling.
Notably, as Marcus recalled in his Nobel lecture, experimental evidence for the inverted-region behavior was not obtained until "almost 25 years after it was predicted" but in that interim "experimentally in some reactions 100% formation of an electronically excited state of a reaction product has been observed" [21]. Further, in a canonical review of ET theory by Barbara, Ratner, and Meyer, it is stated that "there are additional problems with the simple form [of the Marcus-inverted activation barrier]…including an excessively rapid falloff of $k_{\text{ET}}$ with $\Delta G^{\circ}$" [22]. The results of this Letter suggest that such observations—or any experiments in which increasing the driving force leads not to a gradual suppression of the rate but to a sudden transition into an alternative reaction channel—should be revisited for possible reinterpretation. Likewise, cases where the rate saturates and becomes insensitive to changes in driving force may correspond to the extended region of barrierless transition ($\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$) discussed here. Finally, tunable Hamiltonian model systems, where both the driving force and electronic coupling can be systematically varied, may provide a route toward direct experimental testing of these predictions [23, 24, 25, 26, 24].
@@ Acknowledgments
E.A. is grateful to Martin Z. Bazant and Troy Van Voorhis for support, and to Abraham Nitzan and Junghyun Yoon for useful discussions.

### v3

# Adiabatic Electron Transfer in the Barrierless and Marcus-Inverted Regimes

@@ Abstract
Here it is shown that in the adiabatic limit of condensed-phase electron transfer, the onset of barrierless transition occurs at a lower driving force than predicted by the non-adiabatic Marcus formulation. Furthermore, in the adiabatic limit of the Marcus-inverted region, the standard mechanism of electron transfer becomes topologically forbidden. This behavior arises from a topological change in the mapping between the adiabatic and diabatic electronic surfaces, emerging precisely at the onset of the Marcus-inverted region. In this case, alternative mechanisms such as tunneling and non-radiative decay may dominate the rate, typically orders of magnitude slower than the rate calculated from Marcus theory.
@@
but differences can be enormous when $V$ is not much smaller than $\lambda.$ For instance, for aqueous electroreduction of carbon dioxide on a gold electrode in the presence of potassium ions, the reorganization energy has been fitted experimentally by Zhang et al. as $\lambda_{\text{eff}}=0.75$ eV [4], while Qin et al. have measured $\lambda=6.3$ eV in ab initio simulations of the same system [3]. Using these values in Eq. (1) predicts $V=2.1$ eV, which has been corroborated by our own recent ab initio calculation of the coupling value [17]. These results may be critical to quantitatively understand the kinetics of strongly adiabatic electrochemical reactions. Furthermore, in a computational study of organic semiconductors, values of $2V/\lambda>0.5$ have been reported [18], suggesting the possibility that this modification extends beyond electrochemistry to other applications of electron-transfer theory as well.
@@
(iii) $\Delta G^{\circ}<-\lambda$: ET in this inverted region would be forbidden according to a simple Marcus framework and can only occur via alternative routes such as nuclear tunneling or non-radiative transition [19].
We note that the idea that the coupling can suppress the ET rate in the inverted region has appeared in previous work [20, 21, 22]; yet in these treatments a non-adiabatic framework is still assumed and the coupling is treated as a perturbation. However, the large $2V/\lambda$ ratios discussed above are suggestive of a situation that should be viewed primarily in the adiabatic basis, where the Marcus transition state disappears completely.
To justify the proposed behavior in the three cases above, we proceed with the analysis presented in Ref. [16, 23], which represents the Marcus system in the basis of donor (D) and acceptor (A) diabatic states, parametrized by a solvent polarization coordinate $q$. The Hamiltonian is given by
@@
Figure 1: The two level Marcus system under study, where $q$ is the non-dimensionalized reaction coordinate and $E$ is the free energy. The system is shown in (a) the normal region (b) the transition from the normal region to the barrierless region, (c) the barrierless region, and (d) the inverted region in which Marcus theory would predict no transition state.
@@
$k_{\text{ET}}\approx\nu\exp\!\left\{-\frac{\alpha V^{2}}{\lambda}-\frac{(\lambda+\Delta G^{\circ})^{2}}{4\lambda k_{\mathrm{B}}T}\right\}$ (10)
where $\nu$ is the classical attempt frequency. The adiabatic limit implies $\alpha V^{2}/\lambda\gg 1$, and so $k_{\text{ET}}\rightarrow 0$ regardless of the value of the Marcus activation factor. These results suggest that in this limit, quantum mechanical tunneling mechanism or radiationless decay processes may dominate the rate, and these are often expected to be comparatively small [19]. A further implication of these results to liquid systems is that the coupling often depends on the distance between donor and acceptor—so if a reaction is suppressed due to high coupling at small distances, it may be more likely to occur at farther distances.
The framework presented here highlights the importance of continuing to question the standard interpretation of the Marcus inverted region [23, 27]. As Marcus recalled in his Nobel lecture, experimental evidence for the inverted-region behavior was not obtained until "almost 25 years after it was predicted" but in that interim "experimentally in some reactions 100% formation of an electronically excited state of a reaction product has been observed" [28]. The possibility that the general difficulty observing the Marcus inverted region could be due to the adiabaticity of many reactions should be considered, whether results corresponding to the classically forbidden region or he region of extended barrierless transition ($\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$) might explain a given experiment. Tunable Hamiltonian model systems, where both the driving force and electronic coupling can be systematically varied, may provide a route toward direct experimental testing of these predictions [29, 30, 31, 32, 30].
@@ Acknowledgments
E.A. is grateful to Martin Z. Bazant and Troy Van Voorhis for valuable support, and to Abraham Nitzan and Junghyun Yoon for useful discussions.